OpenStudy (anonymous):

Differentiate cot^1 (sqrt(1+x^2) + x) with respect to x?

OpenStudy (anonymous):

is it \[\cot (x) \sqrt{1+x ^{2}}+x \] ?

OpenStudy (anonymous):

Sadly, no :P it's cot inverse.. Or u can say arc cot

OpenStudy (anonymous):

Ohhh i got it.. That's a nice long product rule, chain rule, with trig identity all tied up into one huh?

OpenStudy (anonymous):

And there is no x after cot^-1...

OpenStudy (anonymous):

\[\cot^{-1}(\sqrt{1+x ^{2}}+x) \]

OpenStudy (anonymous):

that right?

OpenStudy (anonymous):

Exactly :)

OpenStudy (anonymous):

Give me a minute to work through it myself. That's a good one. :-p

OpenStudy (anonymous):

Ok man, cheers :)

OpenStudy (anonymous):

\[-((x/\sqrt{x ^{2}+1})+1)/((\sqrt{x ^{2}+1}+x)^{2}+1)\]

OpenStudy (anonymous):

Check that out...Just have to work through it step by step...Hopefully someone else checks that too to make sure I didn't miss something.

OpenStudy (anonymous):

Hmm... So can u explain the steps involved, as in what exactly did u do?

OpenStudy (anonymous):

Okay...Start with the derivative of the inside: \[\sqrt{1+x ^{2}}+x\]

OpenStudy (anonymous):

Can you do that?

OpenStudy (anonymous):

2x/2sqrt(1+x^2). +. 1. ?

OpenStudy (anonymous):

Think about it as (1+x^2)^(1/2) + x

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

you will get x/sqrt(x^x +1) +1 Using the chain rule

OpenStudy (anonymous):

Yeah that's what I've written...

OpenStudy (anonymous):

Okay..and now do you know what the derivative of cot^-1(x)

OpenStudy (anonymous):

-1/1+x^2?

OpenStudy (anonymous):

and it if were a u-substitution it would be -1/1+u^2..with u=inside of the arccot...

OpenStudy (anonymous):

So you end up with (1*x/sqrt(x^x +1) +1) / (u^2 +1) then plug in your u, which was sqrt(x^2 +1) + x

OpenStudy (anonymous):

see it come together?

OpenStudy (anonymous):

Got it @Dio ? I'm about to run.

OpenStudy (anonymous):

Is it x^x or x^2?

OpenStudy (anonymous):

x^2 sorry. the way I typed it the first time. lol

OpenStudy (anonymous):

Trying to get a hang of it. Thanks!

OpenStudy (anonymous):

No problem. That was a tough and tedious one that you just had to watch each piece one by one. You will get it....the more you do, the easier it is to see. :-) good luck.

OpenStudy (anonymous):

Thanks a lot man, u rock :D

OpenStudy (anonymous):

Glad I could help!

OpenStudy (anonymous):

Oh wait.

OpenStudy (anonymous):

It's waaaaay too easy. Don't start differentiating from the beginning. Use some properties of inverse trigonometry to get pi/2 - arccotx, and then differentiate. Thanks for the help anyway, much appreciated.

OpenStudy (anonymous):

Substitution - let x = cot a.